# Autonomous Behavior and Whole-Brain Dynamics Emerge in Embodied Zebrafish Agents with Model-based Intrinsic Motivation

### v1

@@ Appendix D Implementation Details
In all our experiments, we train our agents with PPO using a clipped surrogate objective [47] with $\epsilon^{CLIP}=0.2$ and normalized advantage function computed using Generalized Advantage Estimation (GAE) [52] with $\lambda^{GAE}=0.95$ The policy network is an MLP with two hidden layers [128, 64] optimized by Adam [53] with learning rate $\alpha=0.0003$ and gradients computed over 5 epochs of 1000-step trajectories with a 250-step batch size vectorized across 64 environments. This MLP parameterizes the policy as a 5-dimensional diagonal Gaussian distribution which is sampled to produce actions during training. For evaluation (the experiments in section 4), the policy is deterministic by taking actions as the mean of the distribution. The intrinsic rewards are normalized by dividing by a running estimate of the standard deviation of the sum of discounted rewards with discount factor $\gamma=0.99$, which then supervises the value network MLP with two hidden layers [128, 64] to estimate the expected discounted return using this same discount factor. Both policy and value networks compute on hidden states from separate LSTM blocks (Figure 2A) using a shared embedding from the sensory feature extractors. Image embeddings are obtained from 64x64 pixel observations passed through a three-layer ResNet resulting in outputs with a spatial resolution of 16x16 (closely matching the visual acuity of larval zebrafish at the final layer). Image embeddings were flattened and concatenated with proprioceptive features including joint positions and rotational velocities and their embeddings from a two-layer MLP with hidden size [64, 64].

### v2

@@ Appendix E Implementation Details
In all our experiments, we train our agents with PPO using a clipped surrogate objective [55] with $\epsilon^{CLIP}=0.2$ and normalized advantage function computed using Generalized Advantage Estimation (GAE) [67] with $\lambda^{GAE}=0.95$ . The policy network is an MLP with two hidden layers [128, 64] optimized by Adam [68] with learning rate $\alpha=0.0003$ and gradients computed over 5 epochs of 1000-step trajectories with a 250-step batch size vectorized across 64 environments. This MLP parameterizes the policy as a 5-dimensional diagonal Gaussian distribution which is sampled to produce actions during training. For evaluation (the experiments in section 5), the policy is deterministic by taking actions as the mean of the distribution. Actions take the form of continuous real-valued torques on [1,1][-1,1]. The intrinsic rewards are normalized by dividing by a running estimate of the standard deviation of the sum of discounted rewards with discount factor $\gamma=0.99$, which then supervises the value network MLP with two hidden layers [128, 64] to estimate the expected discounted return using this same discount factor. Both policy and value networks compute on hidden states from separate LSTM blocks (Figure 2A) using a shared embedding from the sensory feature extractors. Image embeddings are obtained from 64x64 pixel observations passed through a three-layer ResNet resulting in outputs with a spatial resolution of 16x16 (closely matching the visual acuity of larval zebrafish at the final layer). Image embeddings were flattened and concatenated with proprioceptive features including joint positions and rotational velocities and their embeddings from a two-layer MLP with hidden size [64, 64].